OpenStudy (anonymous):

Flying against the wind, a jet travels 3080mi in 4 hours. Flying with the wind, the same jet travels 8610mi in 7 hours. What is the rate of the jet in still air and what is the rate of the wind?

OpenStudy (anonymous):

if rate is referring to speed then the initial speed without wind is 770mph compared to 1230mph with wind

OpenStudy (anonymous):

rate is simply reference to something in relation to time, so this is mph of the jet,

OpenStudy (anonymous):

They ask for the Rate of the wind and the Rate of Jet still in Air So there are two answers

OpenStudy (anonymous):

?

OpenStudy (shinalcantara):

let 'x' be the velocity of the jet 'y' be the velocity of the wind flying against the wind the traveled distance is 3080mi in 4hrs note that distance=velocity x time since time is 4 hours and we know that the wind is in opposite direction with the jet then you'll have the equation, x(4) - y(4) = 3080 ---equation 1 flying with the wind you'll have the distance as 8610 mi and time = 7hours applying the same idea you'll have x(7) + y(7) = 8610 ---equation 2 multiplying equation 1 with 7 and equation 2 with 4 you'll have: (4x - 4y = 3080)(7) 28x - 28y = 21,560 ---equation 1' (7x + 7y = 8610)(4) 28x + 28y = 34,440 ---equation 2' subtract equation 1' from 2' 28x + 28y = 34,440 (-) 28x - 28y = 21,560 56y = 12,880 y = 230mph substitute to equation 1 4x - 4y = 3080 4x - 4(230) = 3080 4x = 3080 + 920 4x = 4000 x = 1000mph